OpenStudy (anonymous):

What are the area and perimeter of the sector of the circle shown? C is the center.

OpenStudy (anonymous):

@DelTaVsPi

OpenStudy (anonymous):

Okay so lets find Area first :)

OpenStudy (anonymous):



OpenStudy (anonymous):

Considering that this is a quarter of a circle, we have to take into account that it is important !

OpenStudy (anonymous):

So lets begin!

OpenStudy (anonymous):

sounds good!

OpenStudy (anonymous):

What I would do first is to establish the formula

OpenStudy (anonymous):

find length X width

OpenStudy (anonymous):

which equals 9

OpenStudy (anonymous):

\[A = \frac{\theta }{ 360 }\times \pi \times r^2\]

OpenStudy (anonymous):

no no don't do that, we can't really just time those lengths together

OpenStudy (anonymous):

Okay so, what do you think the radius is to this diagram?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

or. 95

OpenStudy (anonymous):

yes that's right, it's kind of self-explanatory hahaXD

OpenStudy (anonymous):

Anyways...

OpenStudy (anonymous):

How many degrees is this circle?

OpenStudy (anonymous):

360

OpenStudy (anonymous):

wait was r 3?

OpenStudy (anonymous):

Since it is a quarter of a circle The entire circle has 360 degrees, so that means a quarter of a circle is 90 degrees --- Thus, the degree is 90 Degrees :)

OpenStudy (anonymous):

Now we can substitute everything into the formula!

OpenStudy (anonymous):

\[A = \frac{ 90 }{ 360 } \pi \times(3)^2 = 7.068\]

OpenStudy (anonymous):

is is 7.065?

OpenStudy (anonymous):

this is the area, that I've calculated for you:)

OpenStudy (anonymous):

We can now solve the perimeter now!

OpenStudy (anonymous):

The formula of the circumference of the circle is given by: C = \[C = 2\pi \times r\]

OpenStudy (anonymous):

However, we have to still consider that this diagram is this a quarter, sooooo that means that it is a quarter we are dealing with here. Hence. the perimeter of the diagram is given by\[Perimeter = 2\pi \times(3) + 3 + 3 = 10.71\]

OpenStudy (anonymous):

my thing is saying thats wrong...

OpenStudy (anonymous):

Times that by a quarter by the way

OpenStudy (anonymous):

Sorry let me type it up again :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[P = 1/4 x 2\pi \times(3) + 3 + 3 = 10.712\]

OpenStudy (anonymous):

that "x" is times

OpenStudy (anonymous):

awesome! your so smart!

OpenStudy (anonymous):

can you help with one more? ill tag you

OpenStudy (anonymous):

Yeah sure no problem!

OpenStudy (anonymous):

okay i tagged you for my next question

OpenStudy (anonymous):

Also id there is anything you don't understand, please ask me, so that I can explain to you :)

OpenStudy (anonymous):

no it was all right.